OpenStudy (anonymous):

In a class of 30 students, 14 take tea and 20 take coffee. How many take both tea and coffee?

OpenStudy (anonymous):

n(T U C)=n(T)+n(C)-n(T intersection C) 30=14+20-n(T intersection C ) .....find n(T intersection C )